### v2

# Streaming Attention Approximation via Discrepancy Theory

Abstract
Large language models (LLMs) have achieved impressive success, but their high memory requirements present challenges for long-context token generation. In this paper we study the streaming complexity of attention approximation, a key computational primitive underlying token generation.
Our main contribution is BalanceKV, a streaming algorithm for $\epsilon$-approximating attention computations based on geometric process for selecting a balanced collection of Key and Value tokens as per Banaszczyk’s vector balancing theory. We complement our algorithm with space lower bounds for streaming attention computation. Besides strong theoretical guarantees, BalanceKV exhibits empirically validated performance improvements over existing methods, both for attention approximation and end-to-end performance on various long context benchmarks.
1 Introduction
Transformer-based models are the foundation of ongoing artificial intelligence revolution. Their applications span a wide range of domains, from leading-edge language models (LLM) [AAA+23, Tea24a] to text-to-image [RDN+22, Tea23a, Tea22], text-to-video synthesis [Tea24c], coding assistance [Tea23b] and even in multimodal domains across text, audio, image, and video [Ope24]. At the core of these models is the Transformer architecture, powered by the self-attention mechanism [VSP+17], which enables effective capture of pairwise correlations across tokens in an input sequence. As these models scale in size and context length [KMH+20], they face significant computational challenges, particularly in terms of memory usage. Efficiency and accuracy are essential to unlock the full potential of LLMs in generating long sequences.
Space bottlenecks in transformer models.
Most large language models, along with multimodal and video models, adopt an autoregressive, decoder-only architecture. This architecture generates tokens sequentially, applying attention dynamically to each newly generated token. To avoid redundant attention score computations during the generation phase, these models explicitly store the key and value embeddings of previously generated tokens in a cache in each attention layer. Thus, a major challenge is the fact that the memory complexity of storing previously generated key value embeddings scales with both the model size (i.e., the number of layers and attention heads) and, critically, the context size. Additionally, each model session typically requires its own dedicated cache for storing key value embeddings, further exacerbating memory usage. This growing demand has become a significant bottleneck, affecting both memory consumption and computational speed, particularly for models handling long context lengths.
Streaming attention computation.
The main reason for the need of storing the past key and value embeddings is for the attention computation happening inside each self attention layer during token generation after processing a context – to generate the next token, each self attention layer computes the attention between the query embedding of the current token and the key and value embeddings of all the tokens that were previously generated or part of the context. In this paper we study the streaming attention approximation problem – the problem of approximately computing attention using a small amount of space, i.e. without storing all previously seen key and value embeddings. Our main contribution is BalanceKV, a novel provably correct algorithm for streaming attention approximation based on discrepancy theory. The core of our approach is a vector balancing algorithm from discrepancy theory that exploits the geometry of key and value tokens to deduce a small subset of them that well approximates the operations happening inside a self-attention layer. We complement our algorithm with a lower bound on the streaming complexity of approximating attention.
An algorithm for streaming attention approximation can directly be used for compressing the key value cache which stores the past key value embeddings in each layer in an LLM, thus improving the efficiency of LLM token generation. We empirically evaluate BalanceKV both on the problem of approximating attention and on end-to-end generation tasks, showing performance gains.
1.1 Related Work
For discrepancy theory, Banaszczyk’s seminal works [Ban98, Ban12] establishing theoretical guarantees for vector set discrepancy have sparked research in the vector balancing problem [DNTTJ18]. This led to algorithmic developments in both offline [Ban10] and online [BJSS19, ALS21, KRR23] settings. The vector balancing problem has particular relevance to streaming and sublinear algorithms, as minimizing a dataset’s discrepancy yields small subsets that effectively preserve the original dataset’s properties. Recently  [PT20, CKW24] extend these discrepancy theory ideas for kernel density estimation using sublinear memory.
A simple yet effective approach is quantizing previously generated key value embeddings with fewer bits  [YYD+24, YKB+24, DCQW24, KZK+24, LYJ+24, HKM+24, ZYXS24, ZDH24].
Another line of work focuses on token-level pruning, where redundant or less important tokens get evicted from the set of all previously generated key value embeddings [BPC20, ZSZ+24, LDL+24, XTC+23, ZHMK24, LHY+24]. Many of the works in this line have used accumulated attention scores to select important previously generated tokens [ZSZ+24, LHY+24, XTC+23]. Recent works extend those methods to an adaptive way of budget allocation across layer [CZG+24] and head [FCA+24].
1.2 Overview of Our Contributions
In this work we take the token subset selection approach to reduce the memory complexity of LLM token generation: store and maintain only a subset of previously generated key and value embeddings corresponding to a few “important” tokens in the sequence. Of course, the central question is how to define “importance” of tokens. Our approach here is to apply discrepancy theory, which, at a high level, considers a token important if it is crucial to preserving the projection of the total collection of tokens onto some direction in the token space. This leads to the idea of selecting a subset of tokens that is “balanced” simultaneously in every direction. Inspired by the recent breakthrough result of [ALS21] on online discrepancy minimization, we design a method for balancing key-value pairs online using small space, namely our BalanceKV algorithm. Interestingly, this algorithm is online, i.e. the importance of a token is determined only by preceding tokens – in sharp contrast with state of the art heuristics for token selection such as PyramidKV [CZG+24] and SnapKV [LHY+24], whose performance, as we show, our algorithm matches or improves upon. Our contributions are:
  1. 1.
    In Section 3 we propose BalanceKV, an algorithm for recursively compressing the set of previously generated tokens using a geometric correlated sampling process based on discrepancy theory. We show that BalanceKV gives provable guarantees for streaming attention approximation under the bounded $\ell_{2}$ norm assumption (Theorem 3.1). Using tools from communication complexity, we also show a lower bound on the memory complexity of any algorithm for streaming attention approximation in Section 3.
    Section 2 contains the formal problem formulation of streaming attention approximation, its applicability to key value cache compression, as well as a technical overview of the main results and techniques of Section 3.
  2. 2.
    In Section 4 we empirically evaluate our algorithm in various settings. In Section 4.1 we show our approach leads to a lower relative error for single layer attention approximation for open-source LLMs including $\mathtt{Llama}$-$\mathtt{3.1}$-$\mathtt{8B}$-$\mathtt{Instruct}$ [DJP+24] and $\mathtt{Ministral}$-$\mathtt{8B}$-$\mathtt{Instruct}$-$\mathtt{2410}$ [Mis24] as compared to uniformly sampling keys and values in the cache. Section 4.1 we also perform ablation studies to show how various parameters in our algorithm affect the relative error for single layer attention approximation. In Sections 4.2 and 4.3 we perform end to end experiments on various benchmarks such as LongBench [BLZ+23] using models of various sizes such as $\mathtt{Llama}$-$\mathtt{3.1}$-$\mathtt{8B}$-$\mathtt{Instruct}$, $\mathtt{Qwen}$-$\mathtt{2.5}$-$\mathtt{14B}$-$\mathtt{Instruct}$ and $\mathtt{Qwen}$-$\mathtt{2.5}$-$\mathtt{32B}$-$\mathtt{Instruct}$ [YYH+24, Tea24d], and Needle in a Haystack [Kam23]. We show that our provable method for attention approximation when applied to key value cache compression performs better compared to previous existing token subset selection heuristics on end to end tasks. Finally in Section 4.4 we present system efficiency metrics regarding our implementation.
2 Technical Overview
In this section, we first set up the formal problem formulation that we tackle, followed by an overview of our techniques and our main results.
2.1 Streaming Attention Approximation: Formulation and Motivation
Autoregressive Transformers generate tokens one by one and each depends on the previously generated tokens. When Transformers process a sequence of tokens, the attention mechanism operates by computing three types of embeddings for each token at every layer: query, key and value. The query and key capture how different tokens interact, while the value is the actual content to be aggregated. Such interactions are quantified by so-called attention scores, obtained by applying the softmax to the inner product between the query of a given token and the keys of all others. These scores determine how much each previous token’s value contributes to the final output. Once the keys and values are computed for a given token, they do not need to be recomputed when generating subsequent tokens.
Formally, suppose that we have a stream of query, key and value embeddings $(q_{1},k_{1},v_{1}),\ldots,(q_{n},k_{n},v_{n})$, that is the $j$-th token is represented as a triplet of ($q_{j},k_{j},v_{j}$) where $q_{j},k_{j},v_{j}\in\mathbb{R}^{d}$ for all $j\in[n]$. Let $K_{j},V_{j}\in\mathbb{R}^{j\times d}$ be matrices defined by stacking those keys and values in their respective rows.
To compute the following at every step $j$ to generate $j+1$ token, is called the streaming attention problem:
$\text{Attn}(q_{j},K_{j},V_{j}):=\text{softmax}\left(\frac{K_{j}\cdot q_{j}}{\sqrt{d}}\right)^{T}\cdot V_{j}.$ (1)
Keeping all of the key-value pairs in the cache is prohibitively expensive, especially for long sequences. Instead, we opt for approximate computation by sampling a few key-value pairs. Specifically, our goal is to construct an algorithm that at every time step $j$ computes an estimator $z_{j}$ for $\text{Attn}(q_{j},K_{j},V_{j})$ in sublinear in $n$ time and memory. In particular for given precision $\varepsilon>0$, $z_{j}$ should satisfy the following error constraint:
$\|z_{j}-\text{Attn}(q_{j},K_{j},V_{j})\|_{2}\leq\varepsilon\left\|\text{softmax}\left(\frac{K_{j}\cdot q_{j}}{\sqrt{d}}\right)\right\|_{2}\|V_{j}\|_{F}.$ (2)
A sublinear in $n$ time and memory algorithm to compute $z_{j}$ will require knowledge of significantly less key-value pairs than $K_{j},V_{j}$, thus reducing the size of the key value cache needed to store them. This motivates the study of streaming attention approximation, as an algorithm for this can directly be used for key value cache compression during LLM token generation. In the next section we discuss how we will construct such an estimator $z_{j}$ at a high level.
2.2 SoftmaxBalance: Attention Approximation via Discrepancy Theory
We now start with presenting the main ideas of our approach. By the definition of softmax, Equation 1 can be written as
$\text{Attn}(q_{j},K_{j},V_{j})=\frac{1}{Z_{j}}\exp\left(\frac{K_{j}\cdot q_{j}}{\sqrt{d}}\right)^{T}\cdot V_{j},$
where for a matrix $A$ we write $\exp(A)$ to denote entry-wise exponential function to $A$ and
$\ Z_{j}:=\sum_{i\in[j]}\exp(\langle k_{i},q_{j}\rangle/\sqrt{d}).$
Our approach to approximate Attn$(q_{j},K_{j},V_{j})$ consists of two subroutines which approximate:$\bf(1)$ Softmax normalization $Z_{j}=\sum_{i\in[j]}\exp(\langle k_{i},q_{j}\rangle/{\sqrt{d}})$, $\bf(2)$ Matrix-vector product between $V_{j}$ and $\exp(K_{j}\cdot q_{j}/\sqrt{d})$. To understand our main idea, suppose we are at the end of the stream (i.e., $j=n$) and we store all key-value pairs $(k_{1},v_{1}),\ldots,(k_{n},v_{n})$. Then for an arbitrary query $q_{n}$ we aim to approximate the matrix-vector product $\exp(K_{n}\cdot q_{n}/\sqrt{d})^{T}\cdot V_{n}=\sum_{i\in[n]}\exp(\langle k_{i},q_{n}\rangle/\sqrt{d})v_{i}$ by choosing a subset of the rows of $K_{n}$ and $V_{n}$ of size at most $n/2$ which corresponds to a compression rate of $0.5$. Suppose we can design an algorithm which splits the set $C$ of all keys and values into two groups $C^{\prime}$ and $C\backslash C^{\prime}$ so that the matrix-vector product function for any query vector $q_{n}$ is roughly equal over $C^{\prime}$ and $C\backslash C^{\prime}$ that is informally,
$\sum_{\{k,v\}\in C^{\prime}}\exp\left(\frac{\langle k,q_{n}\rangle}{\sqrt{d}}\right)v\approx\sum_{\{k,v\}\in C\backslash C^{\prime}}\exp\left(\frac{\langle k,q_{n}\rangle}{\sqrt{d}}\right)v.$
Then, we are able to approximate the matrix-vector product function with either one of the sums above since informally:
$\sum_{\{k,v\}\in C}\exp\left(\frac{\langle k,q_{n}\rangle}{\sqrt{d}}\right)v\approx 2\sum_{\{k,v\}\in C^{\prime}}\exp\left(\frac{\langle k,q_{n}\rangle}{\sqrt{d}}\right)v.$
Therefore, it would suffice to keep the smaller subset of $C^{\prime}$ and $C\backslash C^{\prime}$ as the desired subset of key value embeddings and discard the rest. If we wanted to compress the key value cache to a smaller size by a factor $2^{T}$ for some $T$, we would recursively compress the selected subset using the same procedure $T-1$ more times.
A similar goal is captured by the vector balancing problem studied extensively in discrepancy theory; given a set of vectors $C=\{k_{1},\dots,k_{n}\}\subset\mathbb{R}^{d}$ with $\|k_{j}\|_{2}\leq 1$ for all $j$, partition them into two groups $C^{\prime},C\setminus C^{\prime}$ such that for any $q\in\mathbb{R}^{d}$ it holds $\sum_{k\in C^{\prime}}\langle k,q\rangle\approx\sum_{k\in C\setminus C^{\prime}}\langle k,q\rangle$ with high probability. The Self-Balancing Walk algorithm [ALS21] is a breakthrough result for the above vector balancing problem. However we need to develop an algorithm for the vector balancing problem with respect to function $\exp(\langle k,\cdot\rangle/\sqrt{d})v$ instead of the inner product function $\langle k,\cdot\rangle$.
Our first contribution is to develop an algorithm for our task, building upon the result from the self-balancing walk [ALS21], which essentially randomly partitions the set of keys and values $C$ into $C^{\prime}$ and $C\setminus C^{\prime}$ such that the following holds with high probability under the assumptions that the norms of the query and key embeddings are bounded,
$\left\|\sum_{\{k,v\}\in C^{\prime}}\exp\left(\frac{\langle k,q_{n}\rangle}{\sqrt{d}}\right)v-\sum_{\{k,v\}\notin C^{\prime}}\exp\left(\frac{\langle k,q_{n}\rangle}{\sqrt{d}}\right)v\right\|_{2}\leq O\left(\log(nd)\right)\cdot\max_{j\in[n]}\|v_{i}\|_{2}.$
We refer to this algorithm as SoftmaxBalance, its formal guarantee is presented in Theorem 3.2 and its pseudocode is presented in Algorithm 2. Theorem 3.2 shows that SoftmaxBalance succeeds to divide $C$ into subsets $C^{\prime}$ and $C\backslash C^{\prime}$ which are balanced with respect to function $\exp(\langle k,\cdot\rangle/\sqrt{d})v$ up to an error which only has logarithmic dependence on the size of $C$. In addition, SoftmaxBalance can accept as input value vectors of arbitrary dimension $s$. Therefore, if instead of the value vectors $v_{1},\ldots,v_{n}\in\mathbb{R}^{d}$ we input the set of scalars $v_{1}=\dots=v_{n}=1$, we will get an algorithm for the vector balancing problem with respect to function $\exp(\langle k,\cdot\rangle/\sqrt{d})$. This implies that we can use SoftmaxBalance to compress the key value cache to even approximate the softmax normalization $\sum_{i\in[n]}\exp(\langle k_{i},q_{n}\rangle/{\sqrt{d}})$. We now discuss how to use SoftmaxBalance for streaming attention approximation, i.e. to use it to compute an estimator $z_{j}$ satisfying Equation 2.
2.3 BalanceKV: Implementing SoftmaxBalance in Streaming
For a sequence of $n$ tokens and a given memory budget of $t\ll n$, we aim to design a procedure which applies SoftmaxBalance to select from $n$ key-value embeddings a set of at most $t$ in the streaming setting and can compute an estimator $z_{j}$ satisfying Equation 2 for all steps $j$ in the stream. In the streaming setting one needs to consider the following aspects. As described in the previous section, one iteration of SoftmaxBalance only allows one to select a $n/2$ sized subset of $n$ key-value embeddings, which is higher than the desired budget of $t$ embeddings. This can be easily mitigated by recursively applying SoftmaxBalance $2^{\log(n/t)}$ times, each time halving the set of key-value embeddings. However, this cannot be implemented in the streaming as we have a limited memory budget of $t$ which prohibits us from storing all key-value embeddings during recursion.
To deal with this, we use the classical merge and reduce technique used in the design of streaming algorithms [BHM+21, MG82, GLPW16].
MergeAndReduce algorithm is a recursive binary tree-based approach that allows one to implement SoftmaxBalance recursively in a streaming setting with the total memory not exceeding $\widetilde{O}(dt)$, where $\widetilde{O}(\cdot)$ supresses polynomial in $\log n$ factors, under the assumption that the norms of queries and keys are bounded. The guarantees of MergeAndReduce are presented in Theorem 3.3, its pseudocode in Algorithm 4 and a visual representation in Figure 2. If the norms of all value embeddings in the stream are the same up to constant factors, that is for all $i,j\in[n]$ $0.5\leq\|v_{i}\|_{2}/\|v_{j}\|_{2}\leq 2$, then the outputs of MergeAndReduce can be used to construct an estimator $z_{j}$ satisfying our attention approximation guarantee of equation Equation 2 with precision $\varepsilon$ for $t=\widetilde{O}(\sqrt{d}/\varepsilon)$. However, the value embeddings may have very different norms.
Our main algorithm BalanceKV (pseudocode in Algorithm 1) deals with this issue by grouping the key-value embeddings in the stream according to the norms of the value embeddings, running a separate instance of MergeAndReduce on each group, and combining the outputs of each instance of MergeAndReduce. BalanceKV constructs a final estimator $z_{j}$ satisfying Equation 2 with precision $\varepsilon$ only using $\widetilde{O}(d\sqrt{d}/\varepsilon)$ memory and $\widetilde{O}(d^{2}/\varepsilon^{2})$ runtime per every step $j$ of the stream, assuming the norms of query and key embeddings are bounded. Existing methods [ZHMK24] subsample keys and values independently in the cache, and thus have a $1/\varepsilon^{2}$ dependence on $\varepsilon$ in total memory. The guarantees of BalanceKV are presented in Theorem 3.1.
Finally using the lower bound on the communication complexity of INDEX, we show a lower bound on the memory complexity of any algorithm for streaming attention approximation in Theorem 3.1.
3 Main Theoretical Results
Our main algorithm for streaming attention approximation is BalanceKV. It takes in as input a stream of $n$ tokens $(q_{1},k_{1},v_{1}),(q_{2},k_{2},v_{2}),\ldots,(q_{n},k_{n},v_{n})$ and at every step of the stream outputs an estimate $z_{j}$ to Attn$(q_{j},K_{j},V_{j})$ (see Equation 1 for the definition of Attn(.)(.)( . )) satisfying Equation 2 with precision $\varepsilon$. Assuming that the $\ell_{2}$ norms of $q_{j},k_{j}$ are at most $r$ for all $j$, BalanceKV uses total space $\widetilde{O}(d\sqrt{d}e^{2r^{2}/\sqrt{d}}\cdot 1/\varepsilon)$ and uses $\widetilde{O}(d^{2}e^{4r^{2}/\sqrt{d}}\cdot 1/\varepsilon^{2})$ runtime at each step $j$ of the stream to output $z_{j}$. Our main theorem is as follows.
Theorem 3.1.
For any $r,\varepsilon>0$, any positive integers $n,d$, any set of tokens $(q_{1},k_{1},v_{1}),(q_{2},k_{2},v_{2}),\ldots,\\ (q_{n},k_{n},v_{n})$ where $q_{j},k_{j},v_{j}\in\mathbb{R}^{d}$ satisfy $\|q_{j}\|_{2},\|k_{j}\|_{2}\leq r$ for all $j$, consider an invocation of BalanceKV with
batch size $\text{batch size }t=\widetilde{O}\left(\sqrt{d}e^{2r^{2}/\sqrt{d}}/\varepsilon\right)\text{ and compression rate }2^{-T}\text{ with }T=\log(n/t).$
Then BalanceKV outputs a vector $z_{j}$ satisfying Equation 2 with probability at least $1-1/\text{poly}(n)$ at every step $j$ of the stream. It uses total memory $\widetilde{O}\left(d\sqrt{d}e^{2r^{2}/\sqrt{d}}/\varepsilon\right)$ across all steps of the stream and runtime of $\widetilde{O}\left(d^{2}e^{4r^{2}/\sqrt{d}}/\varepsilon^{2}\right)$ per step of the stream.
A pseudocode of BalanceKV is described in Algorithm 1. At its core BalanceKV relies on our main discrepancy based algorithm, namely SoftmaxBalance– see Section 3.1 for details on SoftmaxBalance. BalanceKV uses the output of SoftmaxBalance to compute estimates of the numerator and denominator of Attn$(q_{j},K_{j},V_{j})$ and returns the desired attention approximation $z_{j}$ for each streamed index $j$. There are two subtleties, however. First, it is important to bucket tokens in the stream according to the norm of the value vectors – see lines 5 and 6. Second, a direct application of SoftmaxBalance would require too much memory space. To ensure small space usage, we apply a classical streaming technique, namely the MergeAndReduce algorithm on top of SoftmaxBalance to reduce the space consumption. The space reduction achieved by MergeAndReduce is by running a logarithmic number of copies of SoftmaxBalance in a tree-like fashion. More details are introduced in Section 3.2.
To summarize, BalanceKV groups tokens in the stream according to the norms of the corresponding value embeddings, runs a separate instance of MergeAndReduce on each group, and combines the outputs of each instance to construct the final estimate for Attn$(q_{j},K_{j},V_{j})$ at each step $j\in[n]$. Next we present SoftmaxBalance and MergeAndReduce. The full proof of Theorem 3.1 is given in appendix SectionA.1. Finally we state the theorem which provides a lower bound on the memory complexity of any algorithm for streaming attention approximation below, its full proof is provided in appendix Section C.
{restatable}
theoremlowerbound Suppose that $r^{2}\leq d$. Any streaming algorithm which on input ($(\{k_{1},v_{1}\},\ldots,\{k_{n},v_{n}\},q)$, $\|q\|_{2},\|k_{i}\|_{2}\leq r$, outputs $z_{q}$ satisfying Equation 2 with probability 0.999 has space complexity
$\Omega\left(\min\left\{\frac{1}{\varepsilon^{2}},d\exp(2r^{2}/\sqrt{d})\right\}\right).$
Algorithm 1 BalanceKV$((q_{j},k_{j},v_{j})_{j=1}^{n},r,t,T,\varepsilon)$
1:input: stream of $n$ tokens $(q_{j},k_{j},v_{j})$, diameter $r$, batch size $t$, compression rate $2^{-T}$, precision parameter $\varepsilon$.
[…]
3:$v_{\text{max}}\leftarrow 0$
4:repeat
5:    Find an index $i$ such that $2^{i}\geq\|v_{j}\|_{2}\geq 2^{i-1}$
[…]
7:   $v_{\text{max}}\leftarrow\max\left\{\|v_{j}\|_{2},v_{\text{max}}\right\}$
[…]
11:    Send $(k_{j},1)$ as input to MR-Denominator
[…]
13:    output: $z_{j}=\frac{\sum_{l=0}^{T}2^{l}\sum_{\{k,v\}\in V^{l}}\exp\left(\frac{\langle k,q_{j}\rangle}{\sqrt{d}}\right)v}{\sum_{l=0}^{T}2^{l}\sum_{\{k,v\}\in K^{l}}\exp\left(\frac{\langle k,q_{j}\rangle}{\sqrt{d}}\right)}$
14:   $j\leftarrow j+1$
15:until token stream ends
3.1 SoftmaxBalance
We now present our main discrepancy based compression algorithm, SoftmaxBalance. Given a sequence of key value embeddings $C=\{(k_{1},v_{1}),\ldots(k_{n},v_{n})\}$ (with key and value embeddings having possibly different dimensions), the goal of SoftmaxBalance is to produce a partition of $C$ into subsets $C^{\prime},C\setminus C^{\prime}$ such that for any query $q\in\mathbb{R}^{d}$ we have that $\sum_{(k,v)\in C^{\prime}}\exp(\langle k,q\rangle/\sqrt{d})v\approx\sum_{(k,v)\in C\setminus C^{\prime}}\exp(\langle k,q\rangle/\sqrt{d})v$ with high probability. Without loss of generality assume that $|C^{\prime}|\leq|C|/2$, we can then output $2\sum_{(k,v)\in C^{\prime}}\exp(\langle k,q\rangle/\sqrt{d})v$ as an approximation to $\sum_{(k,v)\in C}\exp(\langle k,q\rangle/\sqrt{d})v$, thus achieving a factor 2 compression. Its description is presented in Algorithm 2 below. We note that while SoftmaxBalance takes as input a sequence of key and value embeddings, it can nevertheless be used to compute the softmax normalization: we simply run it on the keys, with the corresponding value vector one-dimensional and all equal to $1$ – see line 11 in BalanceKV, where SoftmaxBalance is called within the corresponding invocation of MergeAndReduce with value vectors as $1$s. It’s guarantees are as follows.
Algorithm 2 $\textsc{SoftmaxBalance}((k_{j},v_{j})_{j},r_{\text{key}},r_{\text{value}},\delta)$
1:input: stream of $\leq n$ key-value embeddings $(k_{j},v_{j})$, radii $r_{\text{key}}$, $r_{\text{value}}$: $\max_{j}\|k_{j}\|_{2}\leq r_{\text{key}}$, $\max_{j}\|v_{j}\|_{2}\leq r_{\text{value}}$, probability of failure $\delta$.
2:$R\leftarrow\exp(r_{\text{key}}^{2}/2\sqrt{d})\cdot r_{\text{value}}$
3:$c\leftarrow 30\log(n/\delta)$
4:Initialize zero vector $\eta\leftarrow\{0\}$
5:for $j$ from 1 and until the end of the stream do
6:   $y\leftarrow\left(\exp(\langle k_{i},k_{j}\rangle/\sqrt{d})\langle v_{i},v_{j}\rangle\right)_{i\in[j]}$
7:   if $\left|y^{T}\eta\right|>c\cdot R^{2}$ then FAIL
8:   $p_{j}\leftarrow\frac{1}{2}-\frac{y^{T}\eta}{2c\cdot R^{2}}$
9:   $\eta_{j}\leftarrow\left\{{\begin{array}[]{ll}+1&\text{ with probability}\,p_{j}\\ -1&\text{o.w.}\end{array}}\right.$
10:   Add a new zero coordinate $\eta_{j+1}\leftarrow 0$
11:end for
12:if |$|\{(k_{i},v_{i}):\eta_{i}=1\}|\leq|\{(k_{i},v_{i}):\eta_{i}=-1\}|$ then
13:   output: $\{(k_{i},v_{i}):\eta_{i}=1\}$
14:else
15:   output: $\{(k_{i},v_{i}):\eta_{i}=-1\}$
16:end if
Theorem 3.2.
Given sets $K=\left\{k_{1},\ldots,k_{n}\right\}\subset\mathbb{R}^{d},V=\left\{v_{1},\ldots,v_{n}\right\}\subset\mathbb{R}^{s}$, and failure probability $\delta>0$, define $C$ to be the dataset of pairs $C=\{(k_{1},v_{1}),\ldots,(k_{n},v_{n})\}$. There exists a randomized algorithm, SoftmaxBalance, which outputs a subset $C^{\prime}\subset C$, $|C^{\prime}|\leq|C|/2$, such that, for any vector $q\in\mathbb{R}^{d}$, with probability at least $1-\delta$,
$\left\|\sum_{\{k,v\}\in C^{\prime}}\exp\left(\frac{\langle k,q\rangle}{\sqrt{d}}\right)v-\sum_{\{k,v\}\notin C^{\prime}}\exp\left(\frac{\langle k,q\rangle}{\sqrt{d}}\right)v\right\|_{2}\leq$
$O\left(\sqrt{s}\cdot\log(ns/\delta)\cdot\exp\left(\frac{\|q\|^{2}_{2}}{2\sqrt{d}}\right)\cdot\exp\left(\max_{j\in[n]}\frac{\left\|k_{j}\right\|^{2}_{2}}{2\sqrt{d}}\right)\cdot\max_{j\in[n]}\|v_{j}\|_{2}\right).$
The runtime of SoftmaxBalance is $O((d+s)n^{2})$ and memory is $O((d+s)n)$.
The proof of the above theorem uses the breakthrough result of [ALS21] for the vector balancing problem, one of the main problems in discrepancy theory. Given a set of vectors $k_{1},\ldots,k_{n}$ the result of [ALS21] produces a subset $C$ of these vectors of at most half the size such that for any vector $q$ we have that $\sum_{k\in C}\langle k,q\rangle\approx\sum_{k\in[n]\setminus C}\langle k,q\rangle$ with high probability. Our main contribution is an algorithm for the vector balancing problem with respect to the function $\exp(\langle k,\cdot\rangle/\sqrt{d})v$ as compared to $\langle k,\cdot\rangle$ in the case of [ALS21]. We defer the proof of Theorem 3.2 to Section A.2.
3.2 MergeAndReduce
As briefly mentioned above in Section 3, MergeAndReduce is a streaming version of SoftmaxBalance. The idea is to partition the stream of tokens into batches of size $t$, apply SoftmaxBalance to the batches to reduce the size of each batch by a constant factor, and then repeat recursively – see Fig. 2 in the appendix.
If we set batch size $t$ to be about $1/\varepsilon$ (see Theorem 3.3 below for the more precise setting), we obtain a streaming algorithm that approximates $\sum_{i=1}^{j}\exp(\langle k_{i},q_{j}\rangle/\sqrt{d})v_{i}$ at any point $j$ in the stream using total space $\widetilde{O}(d\sqrt{d}e^{2r^{2}/\sqrt{d}}/\varepsilon)$ and runtime $\widetilde{O}(d^{2}e^{4r^{2}/\sqrt{d}}/\varepsilon^{2})$ per step, where $r$ is an upper bound on the norms of key and query embeddings.
As before, an important aspect is that MergeAndReduce can handle value embeddings of dimension not necessarily equal to that of key and query embeddings. Thus, when run on scalars $v_{i}=1$ for all $i$, it can also be used to approximate softmax normalization at any point $j$ in the stream. This is the main subroutine used in BalanceKV to approximate Attn$(q_{j},K_{j},V_{j})$. Its pseudocode description is presented in Section A.3.1, and its proof is in Section A.3.2. The formal guarantees are
Theorem 3.3.
For any $r,\varepsilon>0$, any set of tokens $(q_{1},k_{1},v_{1}),\ldots,(q_{n},k_{n},v_{n})$ where $q_{j},k_{j}\in\mathbb{R}^{d}$ satisfy $\|q_{j}\|_{2},\|k_{j}\|_{2}\leq r$, $v_{j}\in\mathbb{R}^{s}$ for $s\leq d$ suppose,
batch size $\text{batch size }t=\widetilde{O}(\sqrt{s}e^{2r^{2}/\sqrt{d}}/{\varepsilon})\text{ and compression rate }2^{-T}\text{ with }T=\log(n/t).$
Then MergeAndReduce on input parameters $t,r,d,s,\varepsilon$, outputs at every step $j$ of the stream subsets of key-value embedding pairs $C^{0},\ldots,C^{T}\subset C:=\{(k_{1},v_{1}),\ldots,(k_{n},v_{n})\}$ such that
$z_{j}:=\sum_{i=0}^{T}2^{i}\sum_{\{k,v\}\in C^{i}}\exp\left(\frac{\langle k,q_{j}\rangle}{\sqrt{d}}\right)v$
satisfies with probability at least $1-1/\text{poly}(n)$
$\left\|\sum_{i=1}^{j}\exp\left(\frac{\langle k_{i},q_{j}\rangle}{\sqrt{d}}\right)v_{i}-z_{j}\right\|_{2}\leq\varepsilon j\cdot e^{-r^{2}/\sqrt{d}}\cdot\max_{i\in[n]}\|v_{i}\|_{2}.$
Total memory of the algorithm is $\widetilde{O}(d\sqrt{s}e^{2r^{2}/\sqrt{d}}/{\varepsilon})$, its $j$-th iteration runtime is $\widetilde{O}(dse^{4r^{2}/\sqrt{d}}/{\varepsilon^{2}})$.
4 Experiments
In this section we now present our experimental results. The full details of all sections as well as the experimental setup and implementation can be found in Appendix B.
4.1 Ablation Studies on Single Layer Attention Approximation
We evaluate the effectiveness of BalanceKV for approximating attention in individual layers of $\mathtt{Llama}$-$\mathtt{3.1}$-$\mathtt{8B}$-$\mathtt{Instruct}$ [DJP+24] and $\mathtt{Ministral}$-$\mathtt{8B}$-$\mathtt{Instruct}$-$\mathtt{2410}$ [Mis24] on the TriviaQA dataset from LongBench [BLZ+23]. Specifically, we examine layers 1, 2, and 5 and compare against independent uniform sampling key and value embeddings.
We provide the full experimental details in Section B.1. For each layer, we approximate attention for recent tokens using a compressed cache that retains a fixed number of initial and recent embeddings, alongside intermediate ones selected via BalanceKV, and measure its relative error against exact attention. We vary the compression rate $2^{-T}\in\{1/2,1/4,1/8,1/16\}$.
As shown in Fig. 1, BalanceKV consistently yields lower relative approximation error than approximating attention by uniform sampling past key value pairs across all settings, empirically validating its advantage as predicted by Theorem 3.1.
Figure 1: Comparison of relative errors across different layers of $\mathtt{Llama}$-$\mathtt{3.1}$-$\mathtt{8B}$-$\mathtt{Instruct}$ (left) and $\mathtt{Ministral}$-$\mathtt{8B}$-$\mathtt{Instruct}$-$\mathtt{2410}$ (right) on TriviaQA dataset.
For a fixed dataset and layer, we also analyzed how the performance and runtime of BalanceKV depend on the batch size and compression rate. More precisely, we repeat the single-layer attention approximation experiment TriviaQA and layers 1 and 15 of $\mathtt{Llama}$-$\mathtt{3.1}$-$\mathtt{8B}$-$\mathtt{Instruct}$, for batch size $\in[64,128,256]$ and compression rate $2^{-T}\in[1/2,1/4,1/8]$. The results are presented in Figure 3. As this experiment suggests, the quality of attention approximation increases as the size of the block doubles, while the runtime becomes slower as also proven theoretically.
4.2 End-to-end Evalution on LongBench
Next we evaluate BalanceKV on LongBench dataset [BLZ+23], which tests long-context understanding across tasks like QA, summarization, few-shot learning, synthetic reasoning, and code completion. Specifically, we test a version of uniform length distribution (LongBench-E). During inference, we compress the key-value cache in the prefill stage using a uniform compression rate of (approximately) 0.25 across all methods, while retaining all streamed embeddings during the decoding phase. We compare against StreamingLLM [XTC+23], SnapKV [LHY+24], PyramidKV [CZG+24], and uniform sampling (see Section 4.1), using their implementations from MInference [JLZ+24]. The evaluation follows the LongBench protocol, using three pre-trained models at different scales including $\mathtt{Llama}$-$\mathtt{3.1}$-$\mathtt{8B}$-$\mathtt{Instruct}$, $\mathtt{Qwen}$-$\mathtt{2.5}$-$\mathtt{14B}$-$\mathtt{Instruct}$ and $\mathtt{Qwen}$-$\mathtt{2.5}$-$\mathtt{32B}$-$\mathtt{Instruct}$. The results are reported in Table 1.
Notably, BalanceKV consistently achieves the best overall performance among compression methods and across all models, demonstrating its effectiveness in preserving model quality for cache compression.
The full experimental setup details can be found in Section B.2.
Method qasper multi hotpotqa 2wiki gov multinews trec triviaqa samsum p.count p.ret lcc repo-p average Qwen2.5-32B-Instruct Exact (Baseline) 44.56 50.65 69.14 60.39 21.3 19.52 75.33 81.14 43.17 22.0 99.67 50.9 35.22 51.77 StreamingLLM 20.12 34.35 51.84 48.23 19.09 17.10 61.00 51.14 28.52 23.33 41.33 39.19 26.54 35.52 PyramidKV 34.47 46.33 67.78 55.92 15.16 15.39 69.33 63.24 40.36 22.67 99.33 48.32 34.35 47.13 SnapKV 36.21 46.78 66.64 57.02 16.35 16.07 70.33 77.53 41.08 22.0 99.33 49.04 35.62 48.77 Uniform 39.28 43.82 64.82 57.84 23.10 19.50 73.00 81.63 39.70 22.00 92.00 44.97 32.19 48.76 BalanceKV 40.14 43.17 64.46 58.06 22.26 20.32 73.00 80.68 41.07 22.33 92.0 44.95 32.43 48.84 Qwen2.5-14B-Instruct Exact (Baseline) 43.39 52.63 64.06 53.71 28.07 22.4 74.67 88.75 44.81 22.33 99.0 63.69 46.3 54.14 StreamingLLM 20.73 32.62 49.93 42.39 21.63 18.69 59.67 74.96 29.57 11.67 63.0 46.16 32.25 38.71 PyramidKV 31.76 46.6 62.83 50.0 19.08 17.68 65.0 85.52 42.61 22.0 99.33 60.62 44.4 49.8 SnapKV 32.95 47.53 61.96 50.34 20.29 18.28 60.67 88.75 42.97 22.67 99.33 61.32 45.84 50.22 Uniform 37.07 41.69 61.11 49.96 29.18 22.67 71.67 87.89 40.34 22.0 84.33 58.0 42.57 49.88 BalanceKV 37.02 41.96 61.74 50.9 29.26 22.64 71.67 88.1 41.14 23.67 87.67 58.64 43.63 50.62 Llama-3.1-8B-Instruct Exact (Baseline) 42.87 48.54 52.05 38.6 31.31 22.07 71.67 91.85 42.36 20.37 98.13 49.62 42.73 50.17 StreamingLLM 20.65 30.71 39.14 32.43 23.10 18.70 58.00 83.87 28.85 20.36 97.26 33.69 30.46 39.79 PyramidKV 33.86 39.75 47.12 35.96 20.03 17.78 63.67 90.76 40.21 20.4 98.97 45.25 39.51 45.64 SnapKV 33.91 42.55 49.09 36.13 20.48 17.67 62.0 91.7 40.23 20.33 98.86 46.7 39.86 46.12 Uniform 37.18 37.15 47.49 37.82 27.56 21.06 68.67 90.48 36.13 20.33 96.26 45.72 37.09 46.38 BalanceKV 35.75 37.04 46.37 36.24 27.09 20.84 69.0 90.88 37.88 20.39 96.65 48.45 41.4 46.77
Table 1: Comparison of various cache compression methods on LongBench-E using various models. The best results among compression methods for each model are highlighted in bold.
4.3 Needle-In-A-Haystack Benchmark
We evaluate BalanceKV to the “Needle-In-A-Haystack“ benchmark [Kam23], comparing it against SnapKV, PyramidKV, StreamingLLM, and uniform sampling using the $\mathtt{Llama}$-$\mathtt{3.1}$-$\mathtt{8B}$-$\mathtt{Instruct}$. The test challenges the model to retrieve a specific sentence (the “needle”) embedded at an arbitrary position within a long context (the “haystack”). Following the setup in [FPN+24], we hide the needle at varying depths, from 0% to 100% of the total context length, across documents ranging from approximately 4K to 100K tokens. As in the previous experiments, all methods are evaluated under a fixed compression ratio of approximately 0.25.
To further enhance performance, we introduce an augmented version of BalanceKV that deterministically preserves a small set of tokens whose key embeddings are strongly anti-correlated with the rest. The standard BalanceKV procedure is then applied to the remaining tokens within each layer. As a result, BalanceKV achieves an average accuracy of 0.99, outperforming SnapKV (0.83), PyramidKV (0.90), StreamingLLM (0.31), and uniform sampling (0.90). Detailed heatmaps of performance across different context lengths and needle depths are in Figure 4 in Section B.3.
4.4 System Efficiency Metrics
Finally, we measure wall-clock times for both the prefill stage (including cache compression) and the decoding stage using a random input of length $16{,}384$ tokens, followed by the generation of $1{,}024$ tokens. Results are averaged over $10$ independent runs, with the minimum runtime reported to enhance robustness. The full results are provided in Table 2 in Section B.4.
All compression methods incur some prefill overhead compared to the uncompressed baseline (Exact). While StreamingLLM achieves the fastest decoding speed, it suffers from significantly lower accuracy (see Table 1). Among the remaining methods, BalanceKV achieves the lowest prefill latency and consistently delivers the best trade-off between efficiency and accuracy. This demonstrates that our discrepancy-based approach not only scales well in theory but also brings practical gains in end-to-end system performance, making it a compelling choice for real-world deployment scenarios.
Appendix A Full proofs of the Algorithm
A.1 Proof of Theorem 3.1
Finally equipped with Theorem 3.2 and Theorem 3.3 we now state the proof of the main Theorem 3.1.
Proof of Theorem 3.1.
[…]
Performance of the algorithm. At time step $j$, procedure MR-Denominator returns subsets $K_{1},\ldots,K_{T}$ such that
$\left|\sum_{i\in[j]}\exp\left(\frac{\langle k_{i},q_{j}\rangle}{\sqrt{d}}\right)-\sum_{l=0}^{T}2^{l}\sum_{\{k,v\}\in K^{l}}\exp\left(\frac{\langle k,q_{j}\rangle}{\sqrt{d}}\right)\right|\leq\varepsilon\cdot j\cdot e^{\frac{-r^{2}}{\sqrt{d}}}\leq\varepsilon\cdot\sum_{i\in[j]}\exp\left(\frac{\langle k_{i},q_{j}\rangle}{\sqrt{d}}\right).$
[…]
$\left\|\sum_{\{k,v\}\in P_{i,j}}\exp\left(\frac{\langle k,q_{j}\rangle}{\sqrt{d}}\right)v-\sum_{l=0}^{T}2^{l}\sum_{\{k,v\}\in V_{i}^{l}}\exp\left(\frac{\langle k,q_{j}\rangle}{\sqrt{d}}\right)v\right\|_{2}$
$\leq\varepsilon|P_{i,j}|\cdot e^{-r^{2}/\sqrt{d}}\cdot 2^{i-1}\leq\frac{\varepsilon|P_{i,j}|\cdot 2^{i-1}}{\sqrt{j}}\cdot\left\|\exp\left(\frac{K_{j}\cdot q_{j}}{\sqrt{d}}\right)\right\|_{2}.$
Now, observe that $\|V_{j}\|_{F}\leq\sqrt{\sum_{i}|P_{i,j}|\cdot 2^{2(i-1)}}$, $j=\sum_{i}|P_{i,j}|$. By the Cauchy-Schwartz inequality,
$\sum_{i}|P_{i,j}|\cdot 2^{i-1}\leq\sqrt{\sum_{i}|P_{i,j}|\cdot 2^{2(i-1)}}\cdot\sqrt{\sum_{i}|P_{i,j}|}.$
[…]
$\frac{\varepsilon}{\sqrt{j}}\cdot\left\|\exp\left(\frac{K_{j}\cdot q_{j}}{\sqrt{d}}\right)\right\|_{2}\sum_{i}|P_{i,j}|\cdot 2^{i-1}\leq\varepsilon\cdot\left\|\exp\left(\frac{K_{j}\cdot q_{j}}{\sqrt{d}}\right)\right\|_{2}\|V_{j}\|_{F},$
therefore we get
$\left\|\frac{\displaystyle\sum_{l=0}^{T}2^{l}\sum_{\{k,v\}\in V^{l}}\exp\left(\frac{\langle k,q_{j}\rangle}{\sqrt{d}}\right)v}{\displaystyle\sum_{l=0}^{T}2^{l}\sum_{\{k,v\}\in K^{l}}\exp\left(\frac{\langle k,q_{j}\rangle}{\sqrt{d}}\right)}-\text{Attn}(q_{j},K_{j},V_{j})\right\|_{2}\leq\frac{2\varepsilon}{1-\varepsilon}\left\|\text{softmax}\left(\frac{K_{j}\cdot q_{j}}{\sqrt{d}}\right)\right\|_{2}\left\|V_{j}\right\|_{F}.$
By rescaling $\varepsilon\to\varepsilon/3$, we get the desired approximation Equation 2. ∎
A.2 Theoretical guarantees of SoftmaxBalance
Algorithm Self-Balancing Walk introduced in [ALS21] receives as input vectors $u_{1},\ldots,u_{n}$ and selects signs for them so that, for any direction, the signed sum of the vectors is balanced along that direction with high probability. The following theorem readily follows from theorem 1.1 in [ALS21]:
Theorem A.1 (Theorem 1.1 in [ALS21]).
For any $n,d\in\mathbb{N}$, there exists a randomized algorithm which receives as input a set of vectors $U=\left\{u_{1},\ldots,u_{n}\right\}\in\mathbb{R}^{d}$ and a parameter $\delta>0$. The algorithm outputs a (random) subset $U^{\prime}\subset U$ such that, for any vector $u\in\mathbb{R}^{d}$, with probability at least $1-\delta$,
$\left|\sum_{i\in U^{\prime}}\left\langle u_{i},u\right\rangle-\sum_{i\notin U^{\prime}}\left\langle u_{i},u\right\rangle\right|\leq O\left(\log(n/\delta)\cdot\max_{i\in[n]}\left\|u_{i}\right\|_{2}\cdot\|u\|_{2}\right).$
[…]
1:input: stream of $\leq n$ vectors $u_{j}$, radius $r$: $\max_{j}\|u_{j}\|_{2}\leq r$, probability of failure $\delta$.
2:$c\leftarrow 30\log(n/\delta)$
3:$U_{-},U_{+}\leftarrow\emptyset$
4:for $i$ from 1 and until the end of the stream do
5:   if $\left|\sum_{u\in U_{+}}\langle u,u_{i}\rangle-\sum_{u\in U_{-}}\langle u,u_{i}\rangle\right|>c\cdot r^{2}$ then
6:      Fail
7:   end if
8:   $p_{i}\leftarrow\frac{1}{2}-\frac{\sum_{u\in U_{+}}\langle u,u_{i}\rangle-\sum_{u\in U_{-}}\langle u,u_{i}\rangle}{2c\cdot r^{2}}$
9:   $\varepsilon_{i}\leftarrow+$ with probability $p_{i}$, and $\varepsilon_{i}\leftarrow-$ with probability $1-p_{i}$
10:   $U_{\varepsilon_{i}}\leftarrow U_{\varepsilon_{i}}\cup\{k_{i}\}$
11:end for
12:if $|U_{+}|\leq|U_{-}|$ then
13:   output: $U_{+}$
14:else
15:   output: $U_{-}$
16:end if
Proof of Theorem 3.2.
Define for any $k\in\mathbb{R}^{d}$ an embedding function $\varphi(k):$
$\varphi(k)=\left(\frac{(k/d^{0.25})^{\otimes i}}{\sqrt{i!}}\right)_{i\geq 0}.$
It is easy to see that for any two vectors $k,q\in\mathbb{R}^{d}$
$\langle\varphi(k),\varphi(q)\rangle=\exp\left(\frac{\langle k,q\rangle}{\sqrt{d}}\right),$
and for any $k\in\mathbb{R}^{d}$
$\|\varphi(k)\|^{2}_{2}=\exp\left(\frac{\|k\|^{2}_{2}}{\sqrt{d}}\right).$
Consider the set of vectors $\varphi(k_{1})\otimes v_{1},\ldots,\varphi(k_{n})\otimes v_{n}$. Run the Self-Balancing Walk algorithm on the set of vectors $\varphi(k_{1})\otimes v_{1},\ldots,\varphi(k_{n})\otimes v_{n}$ with failure parameter set to $\delta/s$ and denote by $C^{\prime}$ and $C\backslash C^{\prime}$ the partition of $C$ returned by the algorithm. Observe that, even though vectors $\varphi(k_{i})\otimes v_{i}$ are infinite dimensional, Self-Balancing Walk still can be implemented. The algorithm never has to keep these vectors in the memory because the only operation which requires the knowledge of the embeddings  – the inner product  – can be performed if we just store vector pairs $\{k_{i},v_{i}\}$:
$\langle\varphi(k_{i})\otimes v_{i},\varphi(k_{j})\otimes v_{j}\rangle=\exp\left(\frac{\langle k_{i},k_{j}\rangle}{\sqrt{d}}\right)\cdot\langle v_{i},v_{j}\rangle.$
Denote by $e_{1},\ldots,e_{s}$ the standard orthonormal basis in $\mathbb{R}^{s}$. By Theorem A.1, for any $i\in[s]$ with probability $1-\delta/s$
|$\left|\sum_{\{k,v\}\in C^{\prime}}\langle\varphi(k)\otimes v,\varphi(q)\otimes e_{i}\rangle-\sum_{\{k,v\}\notin C^{\prime}}\langle\varphi(k)\otimes v,\varphi(q)\otimes e_{i}\rangle\right|$ (3)
$\leq O\left(\log(ns/\delta)\cdot\max_{\{k,v\}\in C}\left\|\varphi(k)\otimes v\right\|_{2}\cdot\|\varphi(q)\otimes e_{i}\|_{2}\right),$ (4)
and so with probability at least $1-\delta$ all of the above inequalities hold simultaneously. To simplify the right hand side, notice that $\left\|\varphi(k)\otimes v\right\|_{2}=\exp\left(\frac{\|k\|^{2}_{2}}{2\sqrt{d}}\right)\cdot\|v\|_{2}$ and $\|\varphi(q)\otimes e_{i}\|_{2}=\exp\left(\frac{\|q\|^{2}_{2}}{2\sqrt{d}}\right)$. Observe that for any $i$ $\langle\varphi(k)\otimes v,\varphi(q)\otimes e_{i}\rangle=\langle\varphi(k),\varphi(q)\rangle\cdot[v]_{i}=\exp\left(\frac{\langle k,q\rangle}{\sqrt{d}}\right)\cdot[v]_{i}$ , where by $[v]_{i}$ we denote the $i$-th coordinate of the vector $v$. Therefore, the left hand side of the expression above is simply the absolute value of the $i$-th coordinate of the vector
$\sum_{\{k,v\}\in C^{\prime}}\exp\left(\frac{\langle k,q\rangle}{\sqrt{d}}\right)v-\sum_{\{k,v\}\notin C^{\prime}}\exp\left(\frac{\langle k,q\rangle}{\sqrt{d}}\right)v.$
Thus, Equation 3 provides a uniform upper bound on the absolute values of coordinates of the above vector. Since the $l_{\infty}$ norm of a vector is the maximum of the absolute values of its coordinates,
$\left\|\sum_{\{k,v\}\in C^{\prime}}\exp\left(\frac{\langle k,q\rangle}{\sqrt{d}}\right)v-\sum_{\{k,v\}\notin C^{\prime}}\exp\left(\frac{\langle k,q\rangle}{\sqrt{d}}\right)v\right\|_{\infty}$
$=\max_{i\in[s]}\left|\sum_{\{k,v\}\in C^{\prime}}\langle\varphi(k)\otimes v,\varphi(q)\otimes e_{i}\rangle-\sum_{\{k,v\}\notin C^{\prime}}\langle\varphi(k)\otimes v,\varphi(q)\otimes e_{i}\rangle\right|$
$\leq O\left(\log(ns/\delta)\cdot\max_{\{k,v\}\in C}\left(\exp\left(\frac{\|k\|^{2}_{2}}{2\sqrt{d}}\right)\cdot\|v\|_{2}\right)\cdot\exp\left(\frac{\|q\|^{2}_{2}}{2\sqrt{d}}\right)\right).$
Finally, we go from bounding the $l_{\infty}$ norm a vector to bounding its $l_{2}$ norm:
$\left\|\sum_{\{k,v\}\in C^{\prime}}\exp\left(\frac{\langle k,q\rangle}{\sqrt{d}}\right)v-\sum_{\{k,v\}\notin C^{\prime}}\exp\left(\frac{\langle k,q\rangle}{\sqrt{d}}\right)v\right\|_{2}$
$\leq\sqrt{s}\left\|\sum_{\{k,v\}\in C^{\prime}}\exp\left(\frac{\langle k,q\rangle}{\sqrt{d}}\right)v-\sum_{\{k,v\}\notin C^{\prime}}\exp\left(\frac{\langle k,q\rangle}{\sqrt{d}}\right)v\right\|_{\infty}$
$\leq O\left(\sqrt{s}\cdot\log(ns/\delta)\cdot\max_{\{k,v\}\in C}\left(\exp\left(\frac{\|k\|^{2}_{2}}{2\sqrt{d}}\right)\cdot\|v\|_{2}\right)\cdot\exp\left(\frac{\|q\|^{2}_{2}}{2\sqrt{d}}\right)\right).$
A.3 MergeAndReduce
A.3.1 Pseudocode for MergeAndReduce
The pseudocode for MergeAndReduce is presented in 4.
1:input: stream of $\leq n$ tokens $(k_{j},v_{j})$, batch size $t$, compression rate $2^{-T}$, precision parameter $\varepsilon$.
2:Let SoftmaxBalance be the algorithm as per Theorem 3.2.
3:Initialize $i$-th level subset $C^{i},i=0,\ldots,T,$ to empty
4:repeat
5:   $C^{0}\leftarrow C^{0}\cup\{\{k_{j},v_{j}\}\}$
6:   if $p$ is not a multiple of $t$ then then
7:      output $C^{0},\ldots,C^{T}$
8:      continue
9:   end if
10:   /*Update subsets every $t$ steps*/
11:   $p\leftarrow j/t$, $i\leftarrow 0$
12:   while $p$ is an integer and until $i=T$ do
[…]
14:      $C^{i}\leftarrow\emptyset$
15:      $i\leftarrow i+1$
16:      $p\leftarrow p/2$
17:   end while
18:   output $C^{0},\ldots,C^{T}$
19:until token stream ends
Algorithm 4 MergeAndReduce$((k_{j},v_{j})_{j},t,T,\varepsilon)$
Figure 2: Illustration of the tree structure of MergeAndReduce
A.3.2 Theoretical guarantees of MergeAndReduce
We now present the full proof of Theorem 3.3.
Proof.
Let us first consider the performance of the procedure at time steps which are multiples of $t$. Note that since in the statement of the theorem $T=\log_{2}(n/t)$, condition until in line while is redundant. Observe that at any such $j$-th step the procedure is an online implementation of the following simple offline recursive algorithm on dataset $\{\{k_{1},v_{1}\},\ldots,\{k_{j},v_{j}\}\}$:
  1. 1.
    Set $p=j/t$ and $i=1$. Split the dataset $\{\{k_{1},v_{1}\},\ldots,\{k_{j},v_{j}\}\}$ into batches $B^{0}_{1},\ldots,B^{0}_{p}$ of size $t$.
  2. 2.
    While $p$ is an integer:
    • Run SoftmaxBalance on the batches $B^{i-1}_{1},\ldots,B^{i-1}_{p}$ independently
    • If $p$ is odd, store the output of SoftmaxBalance on $B^{i-1}_{p}$ in $C^{i}$
    • For every $l$, merge the outputs of SoftmaxBalance on $B^{i-1}_{2l-1}$ and $B^{i-1}_{2l}$ into one batch and store them in $B^{i}_{l}$,
    • Update $p\leftarrow\lfloor p/2\rfloor$, $i\leftarrow i+1$. Stop when $p=1$.
Therefore, we will analyze space complexity and performance guarantees of the above offline algorithm.
Probability of success.
Note that our algorithm performs correctly if each of the calls to SoftmaxBalance produces small error on each of the queries $q$ (as in theorem Theorem 3.2). Throughout the stream, we make $O(n/t)$ calls to SoftmaxBalance, and we apply each to at most $n$ queries, so, it is enough to require that that all SoftmaxBalance have failure probability parameter $\delta=1/\text{poly}(n)$.
Space complexity.
Observe that after each iteration of step 2 the number of batches decreases by a factor of two. The maximum batch size is always bounded by $t$. This is because a batch $B_{l}$ at iteration $i$ of step 2 is a union of SoftmaxBalance$(B_{2l-1})$ and SoftmaxBalance$(B_{2l})$ for $B_{2l-1}$ and $B_{2l}$ at iteration $i-1$ of step 2, and SoftmaxBalance reduced the size of the dataset which it has been applied to at least by a factor of 2.
The memory of the procedure is the collection of memory cells $C^{i}$, and $|C^{i}|\leq t$. Since at time step $j$ at most $\log_{2}(p)\leq\log_{2}(n/t)$ memory cells are occupied, the total memory is bounded by $O(dt\log_{2}(n/t))=O(dtT),$ which, using the $\widetilde{O}$ notation, is equal to $\widetilde{O}(d\sqrt{d}e^{2r^{2}/\sqrt{d}}/{\varepsilon})$.
Performance of the algorithm.
Define $B^{i}=\cup_{l}B^{i}_{l}$ – the data points which remained in the batches after $i$ iterations of step 2. By triangle inequality,
$\left\|\sum_{i=1}^{T}2^{i}\sum_{\{k,v\}\in C^{i}}\exp\left(\frac{\langle k,q_{j}\rangle}{\sqrt{d}}\right)v-\sum_{i=1}^{j}\right.\left.\exp\left(\frac{\langle k_{i},q_{j}\rangle}{\sqrt{d}}\right)v_{i}\right\|_{2}$
$\leq\sum_{i=0}^{T-1}\left\|2^{i+1}\sum_{\{k,v\}\in B^{i+1}\cup C^{i+1}}\right.\left.\exp\left(\frac{\langle k,q_{j}\rangle}{\sqrt{d}}\right)v-2^{i}\sum_{\{k,v\}\in B^{i}}\exp\left(\frac{\langle k,q_{j}\rangle}{\sqrt{d}}\right)v\right\|_{2}$
$\leq\sum_{i=0}^{T-1}2^{i}\left\|\sum_{\{k,v\}\in B^{i+1}\cup C^{i+1}}\right.\left.\exp\left(\frac{\langle k,q_{j}\rangle}{\sqrt{d}}\right)v-\sum_{\{k,v\}\in B^{i}\backslash(B^{i+1}\cup C^{i+1})}\exp\left(\frac{\langle k,q_{j}\rangle}{\sqrt{d}}\right)v\right\|_{2}.$
We will refer to the $i$-th summand (starting from 0) on the right hand side as the error produced by the $i+1$-st iteration of step 2. At the $i+1$-st iteration of step 2 we apply SoftmaxBalance to $p/2^{i}$ batches $B^{i}_{1},B^{i}_{2},\ldots$ of size $t$, and we save the outputs of SoftmaxBalance in batches $C^{i+1},B^{i+1}_{1},B^{i+1}_{2},\ldots$. Therefore, by Theorem 3.2 and triangle inequality, the error vector produced by the procedure at the $i+1$-st iteration of step 2 has $l_{2}$ norm bounded by
$O\left(2^{i}\cdot\sqrt{s}\cdot\log(sn)\cdot\left(\frac{p}{2^{i}}\right)\cdot\ e^{r^{2}/\sqrt{d}}\max_{j\in[n]}\|v_{j}\|_{2}\right)=O\left(\sqrt{s}\cdot\log(sn)\cdot\ p\cdot e^{r^{2}/\sqrt{d}}\max_{j\in[n]}\|v_{j}\|_{2}\right),$
since the error parameter $\delta$ of all instances of SoftmaxBalance is set to $1/\text{poly}(n)$. The $l_{2}$ norm of the total error of our procedure is bounded by
$O\left(\sqrt{s}\cdot\log(sn)\cdot T\cdot p\cdot e^{r^{2}/\sqrt{d}}\max_{j\in[n]}\|v_{j}\|_{2}\right).$
By definition, $p=j/t$. In order to ensure that the statement of the theorem is correct, the upper bound on the $l_{2}$ norm of the error vector of the procedure should be less than the desired error $e^{-r^{2}/\sqrt{d}}\cdot\max_{i\in[n]}\|v_{i}\|_{2}$:
$O\left(\sqrt{s}\cdot\log(sn)\cdot T\cdot\frac{j}{t}\cdot e^{r^{2}/\sqrt{d}}\max_{j\in[n]}\|v_{j}\|_{2}\right)\leq\varepsilon j\cdot e^{-r^{2}/\sqrt{d}}\cdot\max_{j\in[n]}\|v_{j}\|_{2}.$
And, since by definition
$t=O\left(\frac{\log^{2}(sn)\cdot\sqrt{s}\cdot e^{2r^{2}/\sqrt{d}}}{\varepsilon}\right),$
the above inequality holds.
Runtime during one time step. At worst, during $j$-th time step the algorithm has to launch SoftmaxBalance $\log_{2}(p)\leq\log_{2}(n/t)=T$ times on batches of size $t$, so the runtime is bounded by $O(dt^{2}T)$. In the $\widetilde{O}$ notation, the runtime is equal to $\widetilde{O}(d^{2}e^{4r^{2}/\sqrt{d}}/\varepsilon^{2}).$
As the final step, we will analyze the performance of the procedure at time steps $j^{\prime}$ which are not multiples of $t$. Define $j_{t}=\lfloor j^{\prime}/t\rfloor\cdot t$. Note that at any such time step the procedure simply saves the triplet $(q_{j^{\prime}},k_{j^{\prime}},v_{j^{\prime}})$ and outputs the sum of the approximation $z_{j_{t}}$ such that
$\left\|\sum_{i=1}^{j_{t}}\exp\left(\frac{\langle k_{i},q_{j}^{\prime}\rangle}{\sqrt{d}}\right)v_{i}-z_{j_{t}}\right\|_{2}\leq\varepsilon j_{t}\cdot e^{-r^{2}/\sqrt{d}}\cdot\max_{i\in[n]}\|v_{i}\|_{2},$
and $\sum_{i=j_{t}+1}^{j^{\prime}}\exp\left(\frac{\langle k_{i},q_{j}^{\prime}\rangle}{\sqrt{d}}\right)v_{i}$. From the above inequality,
$\left\|\sum_{i=1}^{j^{\prime}}\exp\left(\frac{\langle k_{i},q_{j}^{\prime}\rangle}{\sqrt{d}}\right)v_{i}-\left(z_{j_{t}}+\sum_{i=j_{t}+1}^{j^{\prime}}\exp\left(\frac{\langle k_{i},q_{j}^{\prime}\rangle}{\sqrt{d}}\right)v_{i}\right)\right\|_{2}\leq\varepsilon j_{t}\cdot e^{-r^{2}/\sqrt{d}}\cdot\max_{i\in[n]}\|v_{i}\|_{2},$
as desired. ∎
Appendix B Full Experimental Details
Experiments in Section 4.1 and Section 4.3 are performed on a single NVIDIA A100 GPU with 80GB VRAM, and the rest on a single NVIDIA RTX A6000 GPU with 48GB VRAM.
Implementation Detail.
To enhance the practical performance of our algorithm, we implement BalanceKV with parallel operations. Specifically, we consider the cache embeddings of length $n$ and dimension $d$ as a sequence of blocks with length $b$ and reshape them into a tensor of shape $b\times(n/b)\times d$ Then, BalanceKV is applied in parallel to all blocks of length $b$. For cases where $n$ is not divisible by $b$, we pad the embeddings with zeros. After sign assignment to all embeddings in each block (i.e., line 9 in Algorithm 2), it is reshaped to its original length, and we strictly select $n/2$ embeddings, repeating this process for $T$ iterations.
B.1 Ablation Studies on Single Layer Attention Approximation
In this section we re-state with full details the single layer attention approximation experiments presented in Section 4.1.
We empirically evaluate the performance of BalanceKV for approximating a single attention layer, and compare it with independent uniform sampling. We use the pretrained $\mathtt{Llama}$-$\mathtt{3.1}$-$\mathtt{8B}$-$\mathtt{Instruct}$ [DJP+24] and $\mathtt{Ministral}$-$\mathtt{8B}$-$\mathtt{Instruct}$-$\mathtt{2410}$ [Mis24] and TriviaQA dataset from LongBench [BLZ+23], and consider the $1^{st}$,$2^{nd}$ and $5^{th}$ layers of the models for attention approximation.
For given a prompt with length $n$, we store the corresponding query, key, and value embeddings for all layers. Denote a pair of embeddings in some layer by $(q_{1},k_{1},v_{1}),\ldots,(q_{n},k_{n},v_{n})$ and the goal is to approximate the attention $\text{Attn}(q_{j},K_{j},V_{j})$ for the latest $256$ queries, i.e. $j\in[n-256,n]$. Specifically, we keep several first and recent tokens separately and apply BalanceKV to the intermediate row vectors in $K_{j}$. This is motivated by StreamingLLM [XTC+23] as important contexts are likely contained in the first and latest tokens. We retain the first 256 embeddings and the recent ones from $n-256$ to $j$ and our compressed cache contains tokens whose indices are in $[256]\cup S\cup\{n-256,\dots,j\}$ where $S\subseteq[257,n-256]$ can be obtained from BalanceKV. We explore four compression parameters $T\in\{1,2,3,4\}$ which reduces the cache memory by a factor of $2^{-T}$. Let $z_{j}$ be our approximation using BalanceKV plus the recent and first few embeddings at the stream $j\in[n-256,n]$. We compute relative errors $\|z_{j}-\text{Attn}(q_{j},K_{j},V_{j})\|_{F}/\|\text{Attn}(q_{j},K_{j},V_{j})\|_{F}$ for all $j\in[n-256,n]$, batches, heads and input prompts in the dataset. We repeat this with 10 different random seeds and compute their average and standard deviations. We also compare our method to independent uniform sampling, in which we replace the application of BalanceKV with sampling a $2^{-T}$ fraction of key and value embeddings with indices in $[257,n-256]$ uniformly at random. The results are reported in Figure 1.
Next we present the results of the ablation studies described in Section 4.1 which demonstrate how batch size and compression rate affect the relative error in attention approximation for layers 1 and 15 for $\mathtt{Llama}$-$\mathtt{3.1}$-$\mathtt{8B}$-$\mathtt{Instruct}$.
Batch Size 1/2 1/4 1/8
256 0.0603 0.1190 0.1793
128 0.0320 0.0624 0.0922
64 0.0189 0.0349 0.0508
(a) Layer 1 Runtime (s)
1/2 1/4 1/8
256 0.1036 0.1764 0.2655
128 0.1082 0.1833 0.2741
64 0.1137 0.1921 0.2858
(b) Layer 1 Relative Error
1/2 1/4 1/8
256 0.3920 0.4505 0.5096
128 0.3654 0.3951 0.4256
64 0.3592 0.3753 0.3910
(c) Layer 15 Runtime (s)
1/2 1/4 1/8
256 0.1107 0.1935 0.2798
128 0.1121 0.1952 0.2813
64 0.1141 0.1978 0.2845
(d) Layer 15 Relative Error
Figure 3: Runtime and relative error for across different layers and block sizes. In each figure the rows are corresponding to various batch sizes and columns corresponding to various compression rates
B.2 End-to-End Evaluation on LongBench
We now provide the complete experimental details on the end-to-end evaluation in Section 4.2. We benchmark our algorithm on LongBench dataset [BLZ+23], a comprehensive collection of datasets designed to evaluate the long-context understanding capabilities of large language models. Specifically, we test a version of uniform length distribution (LongBench-E). The benchmark consists of various long-text application scenarios, including single-document question-answering, multi-document question-answering, summarization, few-shot learning, synthetic tasks and code completion. We use BalanceKV to compress the key value cache generated in the prefill stage, and maintain all streamed embeddings $(q_{j},k_{j},v_{j})$ during the token decoding/generation stage. This is because the number of generated tokens is much smaller than the input sequence length. We set $b=256$ and $T=2$, achieving a consistent compression rate of $0.25$ across all inputs.
We evaluate our method against several token-level key value cache compression schemes, including
StreamingLLM  [XTC+23], SnapKV [LHY+24], and PyramidKV [CZG+24] as well as uniform sampling described in Section 4.1. We use their implementations from MInference [JLZ+24], and configure their hyperparameters to match a uniform compression rate with $0.25$. We follow the same evaluation metrics from [BLZ+23]. We test them on $\mathtt{Llama}$-$\mathtt{3.1}$-$\mathtt{8B}$-$\mathtt{Instruct}$ as well as a bigger 14B and 32B parameter models $\mathtt{Qwen}$-$\mathtt{2.5}$-$\mathtt{14B}$-$\mathtt{Instruct}$ and $\mathtt{Qwen}$-$\mathtt{2.5}$-$\mathtt{32B}$-$\mathtt{Instruct}$ [YYH+24, Tea24d], with results summarized in Table 1.
Our method consistently achieves the highest average performance among compression methods and across all models, demonstrating its effectiveness in preserving model quality for the cache compression. Notably, on the triviaqa dataset, it achieves near-exact scores compared to uncompressed baselines (e.g., 80.68 vs. 81.14 with Qwen2.5-32B), highlighting its ability to retain high-quality information. We observe that uniform sampling performs competitively with our method and this result justifies that a subset obtained from discrepancy theory has practical impacts on various LLM tasks.
B.3 Needle-In-A-Haystack
In this section we report the plots corresponding to the Needle in a Haystack experiment described in Section 4.3. They are presented in Figure 4.
Figure 4: Comparison of performance on Needle in a Haystack task using $\mathtt{Llama}$-$\mathtt{3.1}$-$\mathtt{8B}$-$\mathtt{Instruct}$. The methods corresponding to figures from top to bottom are StreamingLLM, SnapKV, PyramidKV, Unif. Sampling and BalanceKV respectively.
B.4 System Efficiency Metrics
In this section we present the prefill and decoding time numbers in Table 2 as described in the system efficiency experimental details in Section 4.4.
Method Prefill Time (sec) Decoding Time (sec)
Exact 3.032 37.769
SnapKV 3.755 40.426
PyramidKV 3.748 37.241
StreamingLLM 3.681 40.276
BalanceKV 3.662 38.054
Table 2: Minimum wall-clock runtime (in seconds) over 10 trials for prefill and decoding stages.
Appendix C Lower bound
[…]
The proof will be a reduction to the well-known INDEX problem.
C.1 Reduction to the INDEX problem
Definition C.1 (The INDEX problem).
Alice gets a bit string $x\sim\text{Unif}\{0,1\}^{n}$ and Bob gets $i\sim\text{Unif}[n]$. Then, the goal is to compute $f(x,i)=x_{i}$ on Bob’s end with a single message $m$ from Alice. Denote by $R_{\delta}^{pub,\to}$ the public coin one-way communication complexity of computing a function $f(x,y)$ with error probability at most $\delta$: Alice holds $x$, Bob holds $y$, they share a source of random bits and Alice sends a single message to Bob, after which he must output the correct answer with probability at least $1-\delta$.
Theorem C.2 (Proven in [Woo20]).
$R_{2/3}^{pub,\to}(INDEX)\geq\Omega(n).$
Proof of Theorem 3.1.
Let $c$ be the small constant such that $R_{2/3}^{pub,\to}(INDEX)\geq c\cdot n$.
[…]
The reduction.
Suppose Alice’s input to the INDEX problem is a bit string $x\in\{0,1\}^{n}$ of length $n=\min\{\frac{1}{\varepsilon^{2}},d\exp(2r^{2}/\sqrt{d})\}$. Using public coins, Alice and Bob jointly generate $n/d$ key vectors $\tilde{k}_{1},\ldots,\tilde{k}_{n/d}\sim\text{Unif}\left\{-\frac{r}{\sqrt{d}},\frac{r}{\sqrt{d}}\right\}^{d},$ function $\pi:[n]\to[n/d]\times[d]$ which randomly partitions the $n$ bits into groups of size $d$, and $n$ random signs $\sigma_{1},\ldots,\sigma_{n}\sim\text{Unif}\{-1,1\}$.
Let $\pi(i)_{1}\in[n/d]$ be the first component of $\pi(i)$ and $\pi(i)_{2}\in[d]$ – the second component of $\pi(i)$. Let $e_{1},\ldots,e_{d}$ be the standard orthonormal basis in $\mathbb{R}^{d}$. We build the dataset of $n$ key-value pairs in the following way. We associate with the $i$-th bit the key vector $k_{i}\coloneqq\tilde{k}_{\pi(i)_{1}}$ and the value vector $v_{i}\coloneqq\sigma_{i}\cdot e_{\pi(i)_{2}}$.
Define
$U=\{\{k_{i},v_{i}\}:x_{i}=1\}$
the set of key-value pairs corresponding to entries 1 in Alice’s bit string $x$. Alice instantiates the streaming algorithm for approximating $\text{Attn}(\cdot,K,V)$ with space complexity $\frac{c}{C}\cdot\min\{\frac{1}{\varepsilon^{2}},d\exp(2r^{2}/\sqrt{d})\}$ with a big enough constant $C$ which we specify later and sends the state of it’s memory before reading $q$ to Bob. When Bob receives the message, he uses it to approximate $\text{Attn}(q_{i},K,V)$ where $q_{i}=k_{i}$. If the value written in the only non-zero coordinate of $v_{i}$ is larger than $\frac{1}{40}\cdot\frac{\exp(r^{2}/\sqrt{d})}{\max\{d\exp(r^{2}/\sqrt{d}),|U|\}}$, Bob reports that the $i$-th bit of Alice’s string is equal to 1, and otherwise  - 0.
Analysis of the reduction.
Proof sketch.
Before moving to formal proofs we briefly outline the main idea of the analysis. Observe that, by the choice of key vectors, any $\exp(\langle k_{j},q_{i}\rangle/\sqrt{d})$ for $j\neq i$ is in expectation insignificantly small compared to $\exp(\langle k_{i},q_{i}\rangle/\sqrt{d})$  – sometimes we will even refer to these terms as “noise”. This statement is formalized in Lemma C.4. Therefore, when Bob computes an approximation to $\text{Attn}(q_{i},K,V)$, he will observe a large value in the coordinate where $v_{i}$ is non-zero if $\{k_{i},v_{i}\}\in U$ and a small value otherwise.
Lemma C.3.
$\mathbb{E}_{x,y\sim\text{Unif}\left\{-\frac{r}{\sqrt{d}},\frac{r}{\sqrt{d}}\right\}^{d}}[\exp(C\langle x,y\rangle/\sqrt{d})]=\Theta(1)$ for any constant $C$.
Proof.
$\mathbb{E}_{x,y\sim\text{Unif}\left\{-\frac{r}{\sqrt{d}},\frac{r}{\sqrt{d}}\right\}^{d}}[\exp(C\langle x,y\rangle/\sqrt{d})]$
$=\left(\frac{1}{2}\exp(Cr^{2}/d^{3/2})+\frac{1}{2}\exp(-Cr^{2}/d^{3/2}))\right)^{d}=\text{cosh}\left(\frac{Cr^{2}}{d^{3/2}}\right)^{d},$
$1\leq\exp(C^{2}r^{4}/4d^{2})\leq\text{cosh}\left(\frac{Cr^{2}}{d^{3/2}}\right)^{d}\leq\exp(C^{2}r^{4}/d^{2})\leq\exp(C^{2})$
where we used the assumption that $r^{2}/d\leq 1$. ∎
Lemma C.4.
Fix $i\in[n]$, select $q_{i}=k_{i}$. Let $|U_{i}|$ be the number of key-value pairs in $U\setminus\{k_{i},v_{i}\}$ whose value vector has non-zero entry in the same coordinate as $v_{i}$.
If $\{k_{i},v_{i}\}\in U$ then with probability $>1-\frac{1}{1000}\cdot\frac{|U_{i}|}{|U|}$
|$\left|\sum_{\{k,v\}\in U}\exp\left(\frac{\langle k,q_{i}\rangle}{\sqrt{d}}\right)\langle v,v_{i}\rangle\right|\geq\exp(r^{2}/\sqrt{d})-O\left(\sqrt{|U|}\right).$
Otherwise, with probability $>1-\frac{1}{1000}\cdot\frac{|U_{i}|}{|U|}$
|$\left|\sum_{\{k,v\}\in U}\exp\left(\frac{\langle k,q_{i}\rangle}{\sqrt{d}}\right)\langle v,v_{i}\rangle\right|\leq O\left(\sqrt{|U|}\right).$
Proof.
We prove both statements using Chebyshev’s inequality.
In the first case, i.e. when $\{k_{i},v_{i}\}\in U$, the sum contains the term $\exp(\langle k_{i},q_{i}\rangle/\sqrt{d})=\exp(r^{2}/\sqrt{d})$, and otherwise it does not. It therefore remains to upper bound the absolute value of the sum
$X=\sum_{\begin{subarray}{c}\{k,v\}\in U,\\ \{k,v\}\neq\{k_{i},v_{i}\}\end{subarray}}\sigma_{k}\exp\left(\frac{\langle k,q_{i}\rangle}{\sqrt{d}}\right)\langle v,v_{i}\rangle=\sum_{\begin{subarray}{c}\{k,v\}\in U_{i},\\ \{k,v\}\neq\{k_{i},v_{i}\}\end{subarray}}\sigma_{k}\exp\left(\frac{\langle k,q_{i}\rangle}{\sqrt{d}}\right)\langle v,v_{i}\rangle,$
$\sigma_{k}\sim\text{Unif}\{-1,1\}$, which effectively introduces “noise” in Bob’s estimate of whether $x_{i}=1$. We upper bound this sum now.
$\mathbb{E}[X]=0$, and $Var(X)=|U_{i}|\cdot Var_{x,y\sim\text{Unif}\left\{-\frac{r}{\sqrt{d}},\frac{r}{\sqrt{d}}\right\}^{d}}(\exp(\langle x,y\rangle/\sqrt{d}))\leq O\left(|U_{i}|\right)$ because
$\begin{split}Var_{x,y\sim\text{Unif}\left\{-\frac{r}{\sqrt{d}},\frac{r}{\sqrt{d}}\right\}^{d}}(\exp(\langle x,y\rangle/\sqrt{d}))\leq\mathbb{E}_{x,y\sim\text{Unif}\left\{-\frac{r}{\sqrt{d}},\frac{r}{\sqrt{d}}\right\}^{d}}[\exp(2\langle x,y\rangle/\sqrt{d})]\leq\exp(4),\end{split}$
by Lemma C.3. We therefore get by Chebyshev’s inequality
$\Pr\left[|X|\geq 1000\sqrt{|U|}\right]\leq\frac{1}{1000}\cdot\frac{|U_{i}|}{|U|}.$
Therefore, with probability $1-\frac{1}{1000}\cdot\frac{|U_{i}|}{|U|}$
|$\left|\sum_{\{k,v\}\in U}\exp\left(\frac{\langle k,q_{i}\rangle}{\sqrt{d}}\right)\langle v,v_{i}\rangle\right|\geq\exp(r^{2}/\sqrt{d})-1000\sqrt{|U|}.$
In the second case, the entire sum equals $X=\sum_{k\in U_{i}}\sigma_{k}\exp(\langle k,q_{i}\rangle/\sqrt{d})$. As shown above, $\Pr\left[|X|\geq 1000\sqrt{|U|}\right]\leq\frac{1}{1000}\cdot\frac{|U_{i}|}{|U|}$. Hence, with probability $1-\frac{1}{1000}\cdot\frac{|U_{i}|}{|U|}$
|$\left|\sum_{\{k,v\}\in U}\exp\left(\frac{\langle k,q_{i}\rangle}{\sqrt{d}}\right)\langle v,v_{i}\rangle\right|\leq 1000\sqrt{|U|}.$
Corollary C.5.
Suppose bits $i_{1},\ldots,i_{d}$ form a group  – that is, $\pi(i_{1})_{1}=\pi(i_{2})_{1}=\ldots=\pi(i_{d})_{1}$. Then all $v_{i_{1}},\ldots,v_{i_{d}}$ have different non-zero coordinates, and therefore $\sum_{j=1}^{d}|U_{i_{j}}|\leq|U|$.
Therefore, by the union bound argument, the conclusion of Lemma C.4 holds for all $d$ bits which form one group simultaneously with probability 0.999.
Lemma C.6.
Fix a bit $i$. With probability 0.98 the following holds:
  1. 1.
    The error of the approximating algorithm in the only non-zero coordinate of $v_{i}$ is bounded by
    $O\left(\frac{\varepsilon}{\sqrt{d}}\cdot\|\text{softmax}(K\cdot q)\|_{2}\cdot\|V\|_{F}\right).$
  2. 2.
    If the $i$-th bit is 1 then
    |$\left|\sum_{\{k,v\}\in U}\exp\left(\frac{\langle k,q_{i}\rangle}{\sqrt{d}}\right)\langle v,v_{i}\rangle\right|\geq\exp(r^{2}/\sqrt{d})-O\left(\frac{\sqrt{|U|}}{\sqrt{d}}\right),$
    and
    |$\left|\sum_{\{k,v\}\in U}\exp\left(\frac{\langle k,q_{i}\rangle}{\sqrt{d}}\right)\langle v,v_{i}\rangle\right|\leq O\left(\frac{\sqrt{|U|}}{\sqrt{d}}\right).$
    otherwise.
Proof.
We may think that the process of generating the dataset and the approximating streaming algorithm has the following order: first Alice and Bob jointly generate the partition $\pi$, the key vectors $\tilde{k}_{1},\ldots,\tilde{k}_{n/d}$ and the value vectors $v_{1},\ldots,v_{n}$ all using public randomness. To generate Bob’s input position $i\in[n]$ we generate pair $a\sim\text{Unif}[n/d]$, $b\sim\text{Unif}[d]$ and declare $i=\pi^{-1}(a,b)$. We may assume that $a$ is chosen before the datasets $K$ and $V$ are generated, and $b$ – after.
Before $b$ is drawn, the key vector $k_{i}$ of $i=\pi^{-1}(a,b)$ is already defined, as well as the datasets $K$, $V$ and $U$. Alice can therefore already apply the streaming algorithm to $U$, and Bob can already apply it to $q_{i}=k_{i}$. Therefore, the error vector which the streaming algorithm yields when applied to $q_{i}=k_{i}$ is also defined before $b$ is known.
Clearly, there are no more than $0.0001\cdot d$ coordinates in which the error of approximation exceeds $10000\cdot\frac{\varepsilon}{\sqrt{d}}\cdot\|\text{softmax}(K\cdot q)\|_{2}\cdot\|V\|_{F}.$ Since every value vector has only one non-zero entry, there are no more than $0.0001\cdot d$ coordinates where at least $10000\cdot\frac{U}{d}$ of value vectors from $U$ have non-zero value. We call all coordinates which are in neither of these two groups safe. From the above, at least $99\%$ of the coordinates are safe. Recall that $b\sim\text{Unif}[d]$, and choosing $b$ is equivalent to choosing the coordinate in which $v_{i}$ is non-zero. Therefore, with probability 0.99 over the choice of $b$ the only non-zero coordinate of $v_{i}$ is safe.
At the same time, similarly to Lemma C.4, by Chebyshev inequality, if $U_{i}\subset U$ is the set of all key-value pairs in $U$ whose value vector has the same non-zero coordinate as $v_{i}$ then with probability 0.999 if the $i$-th bit is 1 then
|$\left|\sum_{\{k,v\}\in U}\exp\left(\frac{\langle k,q_{i}\rangle}{\sqrt{d}}\right)\langle v,v_{i}\rangle\right|\geq\exp(r^{2}/\sqrt{d})-O\left(\sqrt{|U_{i}|}\right),$
and
|$\left|\sum_{\{k,v\}\in U}\exp\left(\frac{\langle k,q_{i}\rangle}{\sqrt{d}}\right)\langle v,v_{i}\rangle\right|\leq O\left(\sqrt{|U_{i}|}\right).$
otherwise.
By union bounding over these two events, we get that the statement of the lemma is correct with high constant probability.
Conclusion of the proof. Let $U^{i}\subset U$ be the set of all pairs from $U$ with the same key as $\{k_{i},v_{i}\}$. Since Alice’s string is drawn from $\text{Unif}\{-1,1\}^{n}$, with probability 0.9 $|U^{i}|\geq 0.4\cdot d$. This is because every bit in the same group as $k_{i}$ belongs to $U$ with probability 1/2.
Observe that by Chebyshev inequality, with high probability 0.999, the denominator of $\text{softmax}(K\cdot q_{i})$ lies in range
$\left[|U^{i}|\cdot\exp(r^{2}/\sqrt{d})+\frac{1}{5}\cdot|U|,|U^{i}|\cdot\exp(r^{2}/\sqrt{d})+20\cdot|U|\right],$
This is because every summand in the denominator, except for $\exp(\langle k_{i},q_{i}\rangle/\sqrt{d})$, is distributed as $\exp(\langle x,y\rangle/\sqrt{d})$, $x,y\sim\text{Unif}\left\{-\frac{r}{\sqrt{d}},\frac{r}{\sqrt{d}}\right\}^{d}$, and the expectation and the variance of this distribution, as shown in Lemma C.3, is $\Theta(1)$. This range is contained in $\left[\frac{1}{5}\cdot(\max\{de^{r^{2}/\sqrt{d}},|U|\}),20\cdot(\max\{de^{r^{2}/\sqrt{d}},|U|\})\right]$. We will denote the denominator as $D$.
Similarly, by Chebyshev inequality, with probability 0.999 the numerator of $\text{softmax}(K\cdot q_{i})$ lies in
$\left[\sqrt{|U^{i}|\cdot\exp(2r^{2}/\sqrt{d})+\frac{1}{5}\cdot|U|},\sqrt{|U^{i}|\cdot\exp(2r^{2}/\sqrt{d})+200\cdot|U|}\right]$
which, since $|U|\leq d\exp(2r^{2}/\sqrt{d})$, is bounded by $\sqrt{200}\cdot\sqrt{d}e^{r^{2}/\sqrt{d}}$.
Suppose that the $i$-th bit is 1. Then,
  • When $\frac{1}{\varepsilon}\geq\sqrt{d}e^{r^{2}/\sqrt{d}}$, by selecting $|U|=\frac{c}{C}\cdot de^{2r^{2}/\sqrt{d}}$ for some enough constant $C$ the value written in the only non-zero coordinate of $v_{i}$ is at least $\frac{e^{r^{2}/\sqrt{d}}}{D}-\frac{1}{100}\frac{e^{r^{2}/\sqrt{d}}}{D}$ and at most $\frac{1}{100}\frac{e^{r^{2}/\sqrt{d}}}{D}$ otherwise, as follows from Lemma C.4;
  • When $\frac{1}{\varepsilon}<\sqrt{d}e^{r^{2}/\sqrt{d}}$, by selecting $|U|=\frac{c}{C}\cdot\frac{1}{\varepsilon^{2}}$ for some big enough constant $C$ the value written in the only non-zero coordinate of $v_{i}$ is at least $\frac{e^{r^{2}/\sqrt{d}}}{D}-\frac{1}{100}\cdot\frac{1}{\varepsilon\cdot D\cdot\sqrt{d}}$, and at most $\frac{1}{100}\cdot\frac{1}{\varepsilon\cdot D\cdot\sqrt{d}}$ otherwise, as follows from Lemma C.4.
The error which the approximator can have in the non-zero coordinate of $v_{i}$ is bounded by
$10000\frac{\varepsilon}{\sqrt{d}}\cdot\|\text{softmax}(K\cdot q)\|_{2}\cdot\|V\|_{F}\leq 10000\frac{\varepsilon}{\sqrt{d}}\cdot\frac{\sqrt{20}\cdot\sqrt{d}e^{r^{2}/\sqrt{d}}}{D}\cdot\sqrt{|U|},$
as shown in Lemma C.6. Below, we show that this error is smaller than the gap between $\frac{1}{40}\cdot\frac{e^{r^{2}/\sqrt{d}}}{\max\{de^{r^{2}/\sqrt{d}},|U|\}}$ and the value written in the coordinate, which means that, even though the approximator introduces some error, Bob is still capable to tell whether the $i$-th bit is 1 or 0.
  • When $\frac{1}{\varepsilon}\geq\sqrt{d}e^{r^{2}/\sqrt{d}}$, the gap between the value written in the coordinate and $\frac{1}{40}\frac{e^{r^{2}/\sqrt{d}}}{\max\{de^{r^{2}/\sqrt{d}},|U|\}}$ is at least $\frac{1}{1000}\cdot\frac{e^{r^{2}/\sqrt{d}}}{\max\{de^{r^{2}/\sqrt{d}},|U|\}}$, and the error
    $\frac{10000\varepsilon}{\sqrt{d}}\cdot\frac{\sqrt{20}\cdot\sqrt{d}e^{r^{2}/\sqrt{d}}}{D}\cdot\sqrt{|U|}\leq\frac{\varepsilon}{2000}\cdot\frac{\sqrt{d}\exp(2r^{2}/\sqrt{d})}{\max\{de^{r^{2}/\sqrt{d}},|U|\}}\leq\frac{1}{2000}\cdot\frac{e^{r^{2}/\sqrt{d}}}{\max\{de^{r^{2}/\sqrt{d}},|U|\}}$
    by an appropriate choice of $C$.
  • When $\frac{1}{\varepsilon}<\sqrt{d}e^{r^{2}/\sqrt{d}}$, the gap between the value written in the coordinate and $\frac{1}{40}\frac{e^{r^{2}/\sqrt{d}}}{\max\{de^{r^{2}/\sqrt{d}},|U|\}}$ is at least $\frac{1}{1000}\cdot\frac{e^{r^{2}/\sqrt{d}}}{\max\{de^{r^{2}/\sqrt{d}},|U|\}}$ and the error
    $\frac{10000\varepsilon}{\sqrt{d}}\cdot\frac{\sqrt{20}\cdot\sqrt{d}e^{r^{2}/\sqrt{d}}}{D}\cdot\sqrt{|U|}\leq\frac{\varepsilon}{2000}\cdot\frac{\exp(r^{2}/\sqrt{d})}{\max\{de^{r^{2}/\sqrt{d}},|U|\}\cdot\varepsilon}\leq\frac{1}{2000}\cdot\frac{\exp(r^{2}/\sqrt{d})}{\max\{de^{r^{2}/\sqrt{d}},|U|\}}$
    by an appropriate choice of $C$.